# On the extraction of Alcock-Paczyński signal from voids: a novel approach via reconstruction

### v1

@@ 5.2 Sensitivity to the void size
In previous analyses, all voids found by the VIDE algorithm were included in the VGCF measurements. However, small voids increase systematic errors and are often excluded; for instance, Hamaus et al. (2022) discard voids smaller than three times the mean tracer separation ($\mathrm{mps}$). Small voids can raise the fraction of spurious voids, which suppress the VGCF amplitude and its multipoles because spurious voids have a null quadrupole and weaker monopole (Cousinou et al. 2019). While spurious voids mainly affect small voids, their impact persists even with large voids (Cousinou et al. 2019). Moreover, small voids are more affected by nonlinear effects, complicating RSD removal. This work focuses on the dynamical aspect, particularly whether small voids bias the estimation of the parameter $\varepsilon$.
@@
Figure 5: Top panel: comparison of the values of the AP parameter $\varepsilon$ with its error $\sigma_{\varepsilon}$ (error bars), obtained with the fitting procedure described in Section 4, for the analysis with tracers and voids in redshift-space (purple dots) and the analysis with tracers and voids in reconstructed space (orange triangles). Each dot represents the $\varepsilon$ value as a function of the minimum radius $R_{\mathrm{min}}$ for the voids in that specific subsample, expressed in mean tracer separation (mps) units, used for computing VGCF. Bottom panel: improvement factor II, i.e., the ratio between $\sigma_{\varepsilon}$ estimated in redshift space and reconstructed space.
@@
To do so, we repeat the redshift-space and reconstructed-space analyses described in the previous section, progressively removing voids with radii smaller than a threshold value defined as $R_{\mathrm{min}}=N_{\mathrm{s}}\times r_{\mathrm{mps}}$, where $r_{\mathrm{mps}}=13\ h^{-1}\mathrm{Mpc}$ denotes the mean halo separation. The results of this test are shown in Fig. 5, where we present the best-fit values of $\varepsilon$ (top panel) and the improvement factor $I$ (bottom panel), for redshift (orange) and reconstructed (blue) space. The error bars in the top panels correspond to the $1\sigma$ uncertainties. All quantities are shown as a function of the minimum void radius, $R_{\mathrm{min}}$. No significant bias in $\varepsilon$ is found regardless of void size or space. Uncertainty on $\varepsilon$ grows with $R_{\mathrm{min}}$ due to fewer voids, but remains consistently smaller in reconstructed space. The improvement factor $I$ remains stable at approximately $1.3$, with a peak at $1.35$, exhibiting a slight decrease to $1.25$ when small voids are excluded, for voids larger than $3\,R_{\mathrm{min}}$. These results corroborate the conclusion that Zel’dovich reconstruction is effective even for small-scale voids, indicating that their dynamics remain only mildly nonlinear once an appropriate Gaussian smoothing is applied to the data (Schuster et al. 2023). The fact that the improvement in precision is already detectable with only about 3,000 voids is encouraging. It suggests that applying this method to future datasets, expected to provide significantly larger void samples from denser tracer catalogs, and thus include many small voids, could yield even greater gains (e.g., Hamaus et al. 2022). By eliminating the need for a cut at small radii, reconstruction allows us to fully exploit the statistical power of these upcoming surveys. The best-fit values presented so far were obtained by analyzing the joint likelihood of the 100 mock void catalogs, with uncertainties rescaled to reflect those of a single catalog. However, when analyzing the results on a mock-by-mock basis, an interesting pattern emerges. In a few cases, the best-fit value of $\varepsilon$ significantly deviates from unity when all halos are included in the analysis. In these instances, the bias disappears once voids with sizes smaller than $3\times\mathrm{mps}$ are excluded. Notably, this is the same cut adopted by Hamaus et al. (2022) and other analyses, to obtain unbiased results. In contrast, the reconstructed space analysis performs better even in these extreme cases, consistently yielding $\varepsilon$ values that are compatible with unity, even when small-scale voids are included. These findings demonstrate that reconstruction not only improves the precision of the $\varepsilon$ estimate but also enhances the robustness of the void-based analysis. This effect can be clearly observed in Fig. 14 in Appendix B, which reproduces the same analysis shown in Fig. 5 but performed on a single mock (mock 0) instead of the full set of 100 mocks. Notably, a similar stabilizing effect of reconstruction was reported for BAO analyses (Sarpa et al. 2021), which further strengthens the case for reconstruction as a unifying framework to maximize both the precision and reliability of large-scale structure probes.
@@ 6 Conclusions
In this work, we present a novel approach to exploiting cosmic voids for performing the Alcock–Paczynski test. The key innovation lies in conducting the entire analysis, including void detection and cross-correlation measurements, after displacing tracers from their redshift-space positions to their real-space positions using the Zel’dovich approximation. This strategy enables the inclusion of numerous small-scale voids, which are typically excluded in standard analyses, thereby increasing the sample size and, consequently, enhancing the precision of cosmological parameter estimation.
@@
Importantly, performing the analysis in reconstructed space leads to a $\sim 23$ % improvement in the precision of the $\varepsilon$ estimate compared to the corresponding redshift-space analysis. The same precision is achieved when using the real space catalogue, demonstrating that our methodology extracts the best information attainable for the given simulation, void finder, and VGCF model. Moreover, reconstruction significantly reduces the degeneracy between $\beta$ and $\varepsilon$, which complicates redshift-space analyses.
@@
A key advantage of the approach proposed here is that it enables the use of the full void sample, without the need to discard small voids. We have shown that including progressively smaller voids does not degrade the quality of the analysis; on the contrary, it steadily improves its precision as it allows to cleanly extract additional information.
@@
This work will be especially valuable for enhancing the statistical power of VGCF analyses with the upcoming spectroscopic galaxy catalogs from DESI (DESI Collaboration et al. 2022), Euclid (Euclid Collaboration et al. 2024), Roman (Spergel et al. 2015; Dore et al. 2019), and SPHEREx (Dore et al. 2018) satellite missions. Before applying the method to observational data, further testing will be required to refine its performance in the presence of AP distortions in light-cone geometry, improve the modeling by accounting for potential systematic biases introduced by the void-finding algorithm when performing the AP test (see Radinović et al. 2024), and incorporate survey-specific observational biases and selection effects. Another important step will be to investigate how different reconstruction methods impact void properties and cosmological constraints within this framework, as partly explored for example in Maragliano et al. (2025). The novel methodology presented in this paper strengthens the role of cosmic voids for precision cosmology by improving parameter constraints from the Alcock–Paczynski test through reconstruction. It lays a solid foundation for the VGCF and void-based analyses to maximize their impact in the extraction of constraints from upcoming galaxy surveys.
@@ A.1 Sensitivity to the choice of parameters that regulate the reconstruction
To this end, we analyze the VGCF measured after performing reconstructions with varying values of $R_{\mathrm{s}}$ and $f$. For each reconstruction, a new void catalog is extracted from the reconstructed halo fields, and the VGCF is measured using the correct cosmology to avoid introducing AP distortions.
Each test consists of two steps. First, we focus on the quadrupole moment and perform a $\chi^{2}$ test to assess its consistency with the null hypothesis of zero signal. Next, we carry out a full statistical inference analysis, similar to that presented in Section 5, to estimate the best-fit values of $\varepsilon$ and $\beta$.
@@ A.1.1 Sensitivity to the smoothing scale
The reconstruction algorithm requires smoothing the input density field with a Gaussian filter. The optimal choice of the smoothing radius $R_{\mathrm{s}}$ depends primarily on the sample’s mean tracer number density, intrinsic clustering, and potentially on the specific cosmic structures and clustering analysis considered, so it must be determined on a case by case basis. $R_{\mathrm{s}}$ should be large enough to suppress strong nonlinear effects and ensure that the Zel’dovich approximation holds, but not so large that it removes genuine power and systematically underestimates the amplitude of peculiar velocities.
To assess the sensitivity to the choice of $R_{\mathrm{s}}$, we performed a test similar to the one employed by Nadathur et al. (2019) to optimize $R_{\mathrm{s}}$ in their reconstruction. We ran several reconstructions using the correct value of $f$ and assuming the correct cosmological model, while varying $R_{\mathrm{s}}$ in the range $0,15]h^{-1}$Mpc, which includes the typical nonlinear scale around $10\ h^{-1}$Mpc. After each reconstructions voids were identified using VIDE and the VGCF was evaluated with the DP estimator.
To assess the quality of the RSD removal, we evaluate the $\chi^{2}$ difference between the measured quadrupole and the null signal using the jackknife covariance matrix as follows:
@@
Table 2: Table of the $\chi^{2}$ values of the measured quadrupole $\xi_{2}$ fitted with the theoretically expected zero quadrupole $\xi_{2}^{\mathrm{th}}$ against different smoothing scales $R_{\mathrm{s}}$.
The best-fit values of $\varepsilon$ and $\beta$ are shown in Fig. 7 as a function of $R_{\mathrm{s}}$ and compared to the expected values (horizontal dashed lines). For values of $R_{\mathrm{s}}$ smaller than the optimal one, the reconstruction yields nonphysical negative values of $\beta$ reflecting the spurious negative quadrupole caused by over-correction (bottom panel). Conversely, for $R_{\mathrm{s}}$ values larger than the optimal scale, the quadrupole becomes positive due to residual outflows, and the best-fit analysis returns a positive $\beta$ to match the signal. On the contrary, $\varepsilon$ (top panel) shows a much weaker sensitivity to the choice of smoothing scale. Its best-fit value remains consistent with the expected value within approximately $\sim 1\sigma$ across the full range of $R_{\mathrm{s}}$ values considered.
@@ A.1.2 Sensitivity to the input $\beta$ parameter
The top panels of Fig. 8 show the quadrupole moments computed using different $f$ values in the reconstruction. Different colors, as indicated in the vertical bar, correspond to various choices for $f$. To avoid overcrowding the plot we separate the reconstructions performed using $f<f^{\mathrm{true}}$ (left) from those performed with $f>f^{\mathrm{true}}$ (right).
We again use the quadrupole as a proxy for assessing reconstruction quality, where significant deviations from zero indicate reconstruction inaccuracies. To aid the visual interpretation, we plot the residuals (divided by the error) relative to the reference quadrupole in the bottom panels. Visual inspection reveals no significant systematic departures from the reference reconstruction and no discernible trend with increasing or decreasing $f$. The estimated reduced $\chi^{2}$ values confirm the visual inspection, with most values close to unity and none reaching 2. Finally, in Fig. 9, we present the best-fit values of $\varepsilon$ and the output $\beta$ as a function of the input $f$. The results demonstrate that both parameters are largely insensitive to the choice of input $f$, remaining consistent with their expected values across the tested range.

### v2

@@ 5.2 Sensitivity to the void size
In previous analyses, all voids found by the VIDE algorithm are included in the VGCF measurements. However, small voids increase systematic errors and are often excluded; for instance, Hamaus et al. (2022) discard voids smaller than three times the mean tracer separation ($\mathrm{mps}$). Small voids can raise the fraction of spurious voids, which suppress the VGCF amplitude and its multipoles because spurious voids have a null quadrupole and weaker monopole (Cousinou et al. 2019). While spurious voids mainly affect small voids, their impact persists even with large voids (Cousinou et al. 2019). Moreover, small voids are more affected by nonlinear effects, complicating RSD removal. This work focuses on the dynamical aspect, particularly whether small voids bias the estimation of the parameter $\varepsilon$.
@@
Figure 5: Top Panel: comparison of the values of the AP parameter $\varepsilon$ with its error $\sigma_{\varepsilon}$ (error bars), obtained with the fitting procedure described in Section 4, for the analysis with tracers and voids in redshift-space (orange dots) and the analysis with tracers and voids in reconstructed space (blue triangles). Each dot represents the $\varepsilon$ value as a function of the minimum radius $R_{\mathrm{min}}$ for the voids in that specific subsample, expressed in mean tracer separation (mps) units, used for computing VGCF. Bottom Panel: improvement factor II, i.e., the ratio between $\sigma_{\varepsilon}$ estimated in redshift space and reconstructed space.
@@
To do so, we repeated the redshift-space and reconstructed-space analyses described in the previous section, progressively removing voids with radii smaller than a threshold value defined as $R_{\mathrm{min}}=N_{\mathrm{s}}\times r_{\mathrm{mps}}$, where $r_{\mathrm{mps}}=13\ h^{-1}\mathrm{Mpc}$ denotes the mean halo separation. The results of this test are shown in Fig. 5, where we present the best-fit values of $\varepsilon$ (top panel) and the improvement factor $I$ (bottom panel), for redshift (orange), and reconstructed (blue) space. The error bars in the top panels correspond to the $1\sigma$ uncertainties. All quantities are shown as a function of the minimum void radius, $R_{\mathrm{min}}$. No significant bias in $\varepsilon$ is found regardless of void size or space. Uncertainty on $\varepsilon$ grows with $R_{\mathrm{min}}$ due to fewer voids, but remains consistently smaller in reconstructed space. The improvement factor $I$ remains stable at approximately $1.3$, with a peak at $1.35$, exhibiting a slight decrease to $1.25$ when small voids are excluded, for voids larger than $3\,R_{\mathrm{min}}$. These results corroborate the conclusion that Zel’dovich reconstruction is effective even for small-scale voids, indicating that their dynamics remain only mildly nonlinear once an appropriate Gaussian smoothing is applied to the data (Schuster et al. 2023). The fact that the improvement in precision is already detectable with only about 3,000 voids is encouraging. It suggests that applying this method to future datasets, expected to provide significantly larger void samples from denser tracer catalogs, and thus include many small voids, could yield even greater gains (e.g., Hamaus et al. 2022). By eliminating the need for a cut at small radii, reconstruction allows us to fully exploit the statistical power of these upcoming surveys. The best-fit values presented so far were obtained by analyzing the joint likelihood of the 100 mock void catalogs, with uncertainties rescaled to reflect those of a single catalog. However, when analyzing the results on a mock-by-mock basis, an interesting pattern emerges. In a few cases, the best-fit value of $\varepsilon$ significantly deviates from unity when all halos are included in the analysis. In these instances, the bias disappears once voids with sizes smaller than $3\times\mathrm{mps}$ are excluded. Notably, this is the same cut adopted by Hamaus et al. (2022) and other analyses, to obtain unbiased results. In contrast, the reconstructed space analysis performs better even in these extreme cases, consistently yielding $\varepsilon$ values that are compatible with unity, even when small-scale voids are included. These findings demonstrate that reconstruction not only improves the precision of the $\varepsilon$ estimate but also enhances the robustness of the void-based analysis. This effect can be clearly observed in Fig. 14 in Appendix B, which reproduces the same analysis shown in Fig. 5 but performed on a single mock (mock 0) instead of the full set of 100 mocks. Notably, a similar stabilizing effect of reconstruction was reported for BAO analyses (Sarpa et al. 2021), which further strengthens the case for reconstruction as a unifying framework to maximize both the precision and reliability of large-scale structure probes.
@@ 6 Conclusions
In this work, we present a novel approach to exploiting cosmic voids for performing the Alcock–Paczynski test. The key innovation lies in conducting the entire analysis in reconstructed space, including void detection and cross-correlation measurements, after displacing tracers from their redshift-space positions to their real-space positions using the Zel’dovich approximation. This strategy enables the inclusion of numerous small-scale voids, which are typically excluded in standard analyses, thereby increasing the sample size and, consequently, enhancing the precision of cosmological parameter estimation.
@@
Importantly, performing the analysis in reconstructed space leads to a $\sim 23$ % improvement in the precision of the $\varepsilon$ estimate compared to the corresponding redshift-space analysis. The same precision is achieved when using the real space catalog, demonstrating that our methodology extracts the best information attainable for the given simulation, void finder, and VGCF model. Moreover, reconstruction significantly reduces the degeneracy between $\beta$ and $\varepsilon$, which complicates redshift-space analyses.
@@
Another key advantage of the approach proposed here is that it enables the use of the full void sample, without the need to discard small voids. We have shown that including progressively smaller voids does not degrade the quality of the analysis; on the contrary, it steadily improves its precision as it allows to cleanly extract additional information.
This work presents the first analysis performed entirely in reconstructed space. It marks a conceptual difference from previous studies, which correlated voids identified either in redshift space (e.g., Hamaus et al. 2020, 2022; Verza et al. 2025) or in reconstructed space (e.g., Nadathur et al. 2020; Woodfinden et al. 2022; Radinović et al. 2023) with galaxies in redshift space. The aforementioned analyses provide a broader pool of cosmological information, since they simultaneously exploit RSD and AP effects, constraining both $\beta$ and $\varepsilon$, but at the cost of increased modeling complexity. In particular, the validity of analytical RSD models around voids is still debated, and identifying voids directly in redshift space can introduce systematic effects such as void fragmentation and the violation of pair conservation between real and redshift space, as discussed in Correa et al. (2022). In contrast, our method isolates the AP signal by performing the entire analysis in reconstructed space, where RSDs are largely removed with reconstruction. This disentangling of distortions, as shown by the reduction of the degeneracy between $\beta$ and $\varepsilon$, allows a cleaner and more precise measurement of the AP parameter $\varepsilon$, which directly probes $D_{\mathrm{A}}(z)H(z)$. The strength of our methodology lies in this conceptual simplification: by avoiding explicit RSD modeling and the associated uncertainties, it provides a robust framework for more precise constraint of the AP test with voids.
This methodology, however, also has its own sources of systematics. The reconstruction technique adopted here is sensitive to the choice of the fiducial cosmology and to the physical scale used to smooth the mass-density field traced by the halos. An incorrect choice of these quantities introduces systematic errors in the reconstructed field and, ultimately, in the void identification and analysis. To assess the magnitude of these potential effects, we performed a series of tests using a suite of mock halo catalogs (see Appendix A). These tests show that, while the reconstruction is robust to the choice of the background cosmological model, provided its parameters lie within a plausible range consistent with current observations, it is more sensitive to the choice of the smoothing filter. The latter must therefore be optimized using dedicated, realistic mock catalogs that reproduce the correct number density of tracers. This is particularly important when analyzing real data, where the spatial number density of tracers varies across the survey volume and systematic errors must remain smaller than the already small statistical uncertainties. Importantly, one may also note that the parameter $\beta$, being effectively treated as a nuisance parameter in the post-reconstruction analysis, should naturally absorb any residual reconstruction errors, further mitigating their impact.
In addition to reconstruction-related uncertainties, it is also worth acknowledging that further systematics may arise from the choice of the void-finding algorithm itself. In particular, ZOBOV-based void catalogs (and generically, for some of the mentioned effects, any void catalog from survey data), typically used in observational analyses, may be affected by survey geometry, tracer selection effects, the number density of the sample, and the watershed segmentation procedure, all of which can impact void properties and, potentially, the AP signal extracted from the VGCF. Since our assessment is conducted on cubic simulation boxes, these algorithm-dependent effects are not entirely captured here and must be tested explicitly under realistic observational conditions. More generally, it is worth distinguishing between the systematics that have been explicitly studied in this work, namely those associated with reconstruction, and those that remain to be explored, including void-finder dependent effects.
However, even the systematics already characterized here will need to be reassessed in a more realistic setting, where the AP effect is properly accounted for. The study of these systematics in the context of the era of big data for voids will be the subject of future work, where we plan to investigate them under controlled conditions. Specifically, we will employ lightcone mocks with varying cosmologies, differing from that of the simulations to introduce a realistic AP effect, in order to rigorously quantify and control systematic biases.
In addition, the sensitivity of our analysis to the numerical resolution of the simulations and to the discrete sampling of galaxies represents another important aspect that deserves dedicated investigation in future work, ideally relying on dense light-cones reproducing surveys’ features. By outlining this roadmap, our work establishes a solid foundation for the robust application of reconstructed void-based AP analyses to forthcoming large-scale structure surveys, thereby maximizing the cosmological information extractable from future data.
@@
This work will be especially valuable for enhancing the statistical power of VGCF analyses with the upcoming spectroscopic galaxy catalogs from DESI (DESI Collaboration et al. 2022), Euclid (Euclid Collaboration et al. 2025), Roman (Spergel et al. 2015; Dore et al. 2019), and SPHEREx (Doré et al. 2014). Before applying the method to observational data, further testing will be required to refine its performance in the presence of AP distortions in light-cone geometry, improve the modeling by accounting for potential systematic biases introduced by the void-finding algorithm when performing the AP test (see Radinović et al. 2024), and incorporate survey-specific observational biases and selection effects. Another important step will be to investigate how different reconstruction methods impact void properties and cosmological constraints within this framework, as partly explored for example in Maragliano et al. (2025).
In particular, an interesting direction for future work is to analyze the different effects of reconstruction on distinct void populations, such as void-in-void and void-in-cloud types, which may exhibit different sensitivities to the reconstruction process. More generally, beyond reconstruction alone, it is important to investigate how different void classes contribute to the cosmological signal, as their distinct environments and dynamical properties may encode complementary information.
The novel methodology presented in this paper strengthens the role of cosmic voids for precision cosmology by improving parameter constraints from the Alcock–Paczynski test through reconstruction. It lays a solid foundation for the VGCF and void-based analyses to maximize their impact in the extraction of constraints from upcoming galaxy surveys.
@@ A.1 Sensitivity to the choice of parameters that regulate the reconstruction
To this end, we analyzed the VGCF measured after performing reconstructions with varying values of $R_{\mathrm{s}}$ and $f$. For each reconstruction, a new void catalog is extracted from the reconstructed halo fields, and the VGCF is measured using the correct cosmology to avoid introducing AP distortions. Each test consists of two steps. First, we focused on the quadrupole moment and perform a $\chi^{2}$ test to assess its consistency with the null hypothesis of zero signal. Next, we carried out a full statistical inference analysis, similar to that presented in Section 5, to estimate the best-fit values of $\varepsilon$ and $\beta$.
@@ A.1.1 Sensitivity to the smoothing scale
The reconstruction algorithm requires smoothing the input density field with a Gaussian filter. The optimal choice of the smoothing radius $R_{\mathrm{s}}$ depends primarily on the sample’s mean tracer number density, intrinsic clustering, and potentially on the specific cosmic structures and clustering analysis considered, so it must be determined on a case-by-case basis. $R_{\mathrm{s}}$ should be large enough to suppress strong nonlinear effects and ensure that the Zel’dovich approximation holds, but not so large that it removes genuine power and systematically underestimates the amplitude of peculiar velocities.
To assess the sensitivity to the choice of $R_{\mathrm{s}}$, we performed a test similar to the one employed by Nadathur et al. (2019) to optimize $R_{\mathrm{s}}$ in their reconstruction. We ran several reconstructions using the correct value of $f$ and assuming the correct cosmological model, while varying $R_{\mathrm{s}}$ in the range $0,15]h^{-1}$Mpc, which includes the typical nonlinear scale around $10\ h^{-1}$Mpc. After each reconstructions, voids were identified using VIDE and the VGCF was evaluated with the DP estimator.
To assess the quality of the RSD removal, we evaluated the $\chi^{2}$ difference between the measured quadrupole and the null signal using the jackknife covariance matrix as follows
@@
Table 2: $\chi^{2}$ values. [^6] Table of the $\chi^{2}$ values of the measured quadrupole $\xi_{2}$ fitted with the theoretically expected zero quadrupole $\xi_{2}^{\mathrm{th}}$ against different smoothing scales $R_{\mathrm{s}}$.
@@
The best-fit values of $\varepsilon$ and $\beta$ are shown in Fig. 7 as a function of $R_{\mathrm{s}}$ and compared to the expected values (horizontal dashed lines). For values of $R_{\mathrm{s}}$ smaller than the optimal one, the reconstruction yields unphysical negative values of $\beta$ reflecting the spurious negative quadrupole caused by over-correction (bottom panel). Conversely, for $R_{\mathrm{s}}$ values larger than the optimal scale, the quadrupole becomes positive due to residual outflows, and the best-fit analysis returns a positive $\beta$ to match the signal. On the contrary, $\varepsilon$ (top panel) shows a much weaker sensitivity to the choice of smoothing scale. Its best-fit value remains consistent with the expected value within approximately $\sim 1\sigma$ across the full range of $R_{\mathrm{s}}$ values considered.
@@ A.1.2 Sensitivity to the input $\beta$ parameter
The top panels of Fig. 8 show the quadrupole moments computed using different $f$ values in the reconstruction. Different colors, as indicated in the vertical bar, correspond to various choices for $f$. To avoid overcrowding the plot, we separate the reconstructions performed using $f<f^{\mathrm{true}}$ (left) from those performed with $f>f^{\mathrm{true}}$ (right).
We again used the quadrupole as a proxy for assessing reconstruction quality, where significant deviations from zero indicate reconstruction inaccuracies. To aid the visual interpretation, we plot the residuals (divided by the error) relative to the reference quadrupole in the bottom panels. Visual inspection reveals no significant systematic departures from the reference reconstruction and no discernible trend with increasing or decreasing $f$. The estimated reduced $\chi^{2}$ values confirm the visual inspection, with most values close to unity and none reaching 2. Finally, in Fig. 9, we present the best-fit values of $\varepsilon$ and the output $\beta$ as a function of the input $f$. The results demonstrate that both parameters are largely insensitive to the choice of input $f$, remaining consistent with their expected values across the tested range.